aeroplane flying horizontally 1km above the ground is observed at an elevation of 60o .If after 10 seconds the elevation is observed to be 30o .The uniform speed for an hour of the aeroplane is

Grade:12

1 Answers

Arun
5 years ago
When the elevation was 60, distance of the point of observation to the point on the ground
below the plane = 1xcot60 = 1/sqrt(3) km
 
Similarly when the angle becomes 30, distance becomes = 1xcot30 = sqrt(3) km
 
So, distance travelled = sqrt 3 -1/sqrt 3 km = 2/sqrt3 km
 
time taken = 10sec = 10/3600 hr
 
hence speed = distance / time = 2x3600/10x sqrt(3) = 720/sqrt (3) km/hr = 415.7 km/hr